Exdifferences−version1.1to2.0 Thissheetsummarizesthedifferencesbetweentheoldversion1.1ofexandthenewversion 2.0.ThenewexisavailableasthestandardexontheVAXonthe5thfloorofEvans,andasa newandexperimentalversionin/usr/newontheCory11/70.Itwillsoonbeavailablein /usr/newontheComputerCenterandIngresMachines.SendproblemsovertheBerkeleynetworkto‘‘vax:bill’’. Changestoexistingfeatures Options. Theoptionseditany,edited,fork,hush,printallandstickyhavebeendeletedbecauseoflack ofuse.Thenotifyoptionhasbeenrenamedreport. Thehomeoptionwillsoonbesupersededbytheenvironmentfeatureofversion7UNIXand hasbeendeleted.Similarlythemodeoptionissupersededbytheumaskofversion7andhasalso beendeleted. Thevisualmessageoptionhasbeendeleted;use‘‘mesgn’’atthesystemcommandlevelto inhibitinterconsolemessages. Theiuloptionisreplacedbyamoregeneralmechanismwhichallowsportionsofthebuffer tobeprocessedthroughspecifiedcommands;youcangetiulprocessingonlines1to100ofafile bydoing‘‘1,100!iul’’.Thisreplacesthelines1to100bytheoutputofaniulcommand,givingthe commandtheselinesasinput. Invocation Theoptions−o,−nand−phavebeendeleted. Filenameformation Thealternatefilenameisnowrepresentedas‘#’ratherthan‘` ’,since‘` ’isashell metacharacter.Theeditornowusesashelltoexpandfilenamescontainingshellmetacharacters. Ifyouusecsh,thenyoucanusealltheshellmetasyntaxinformingnewfilenames,includinghome directoryreferenceswith‘˜ ’andvariablesyoudefinein.cshrcusing‘$’. Characterrepresentation Controlcharactersarenowrepresentedas‘ˆ x’;thusacontrolXisprintedas‘ˆ X’;thedelete characterisrepresented‘ˆ ?’. Commandchanges Therehavebeenmajorchangestoopen/visual(incompatibleonesaredescribedbelow). Itisnolongerpossibletodiscardchangesbyrepeatingthequitcommandtwice.Youmust usethevariantformquit!togetoutoftheeditordiscardingchanges.Similarlythevariantforms e!andnext!mustbeusedtoeditanewfileorthenextfilewithoutsavingchangesyouhave made.† Anewformofthe‘!’shellescapereplacestheexpandandtabulatecommands.Thusthe command‘‘1,10expand’’oftheoldversionisreplacedby‘‘1,10!expand’’inthenew.Notealsothat thecommandabbreviationtanolongerreferstothetabulatecommand,whichhasbeendeleted, butratherreferstothenewtagcommand. Theformatoftheargscommandhasbeenchanged;thefilesarenolongernumbered,rather theentireargumentlistisalwaysprintedwiththecurrentfilenameenclosedby‘[’and‘]’. †Lessusefularerewind!andrecover!.

-2-Theformatofthefilecommandoutputhasbeenchanged;theeditorsays‘[Notedited]’in therarecasethatthisistrueratherthansaying‘[Edited]’.Thecommandalsogivesthepercentageofthewayintothebufferthatthecurrentlineis. Theformatofthesetcommandhasbeenimproved;‘‘setall’’nowprintsinathreecolumn format.Thecommands‘‘set%’’,‘‘set!’’and‘‘set` ’’havebeendeleted.Thecommand‘‘set’’now printsinaonelineformatratherthandownthescreen. Thecommandsecho,expand,help,reset,sync,tabulateandxpandhavebeendeleted. Changestoopenandvisual Alargenumberofchangeshavebeenmadetoopenandvisual;wesummarizeonlythemost noticeableoneshere.Seetheattachedreferencecardformoreinformation,and(evenifyouknow howtousevisualalready)youshouldlookatAnIntroductiontoTextEditingwithVi.Wedonot discussanyofthenewcommandsinvisualhere.† Thedeletelinecommandisnowddratherthan\\(\\nolongerworks!.)Infact,dand otheroperatorscannowoperateonlines;thusdLdeletestothelastlineonthescreen.Theshift commands<and>arenowoperators,thus<<and>>nowhavetheeffectthat<and>used tohave. Thecommandvhasbeendeleted;onlyitssynonymzremains.TheKoperationhasbeen movedtom;Khasnomeaninginthenewversion.Theˆ Soperationhasbeendeleted,butˆ G doesasync,andalsoprintssomeinformation.Theˆ Woperationhasbeendeleted(useB).The #,@andˆ Xoperationshavebeendeleted.Todeletetothebeginningofthelineused0;the commandsandxandXaresimilarto#. Duringinputs,ˆ WbacksuplikebratherthanB. Terminalsupporthasbeenvastlyimproved;theeditorwillnowdrivemostanydisplayterminal,usingallterminalfeaturessuchascursoraddressing,cleartoendofline,insertanddelete lineandinsertanddeletecharacter.Tohelpperformanceonslowterminalssomeoptionsarenow setbasedontheintelligenceandspeedoftheterminal;inparticular,thedefaultwindowsizeis 1/2afullscreenat300baud,or2/3ofafullscreenat1200baud. †Itisnowpossibletoeditwiththefocusoftheeditingbeingvisualusingacommandviratherthan exonthecommandline,andusinganew:commandfromwithinvisualtoruncommandmodecommands.